# On the Sharp Input-Output Analysis of Nonlinear Systems under Adversarial Attacks

### v1

@@ Appendix B Preliminaries on sub-Gaussian variables
$\mathbb{E}\left[|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|\right]\geq\Theta\Bigr{(}\frac{\lambda^{4}}{(\sqrt{M}\tau L_{\phi}\sigma_{u})^{3}}\Bigr{)}=\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}.$ (46)
Now, note that due to Lemma B.2, the sub-Gaussian norm of $|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|-\mathbb{E}\left[|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|\right]$ is $\Theta(\sqrt{M}\tau L_{\phi}\sigma_{u})$ and its mean is indeed zero. Then, using (29b), we have
$\mathbb{P}\biggr{(}|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|-\mathbb{E}\left[|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|\right]>-\frac{1}{2}\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}\biggr{)}\geq 1-\exp\Bigr{(}-\Theta\Bigr{(}\frac{1}{\nu^{8}}\Bigr{)}\Bigr{)}.$
Applying (46) to the above inequality incurs
$\mathbb{P}\biggr{(}|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|>\frac{1}{2}\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}\biggr{)}\geq 1-\exp\Bigr{(}-\Theta\Bigr{(}\frac{1}{\nu^{8}}\Bigr{)}\Bigr{)}.$
for all $t$.
Since $1/\nu^{8}$ is greater than 00, the quantity on the right-hand side is a constant strictly greater than 00. We let
$\alpha:=1-\exp\Bigr{(}-\Theta\Bigr{(}\frac{1}{\nu^{8}}\Bigr{)}\Bigr{)}.$ (47)
Since $\bm{U}_{t}^{(\tau)}$ includes $u_{t},\dots,u_{t-\tau}$, any two vectors $\bm{U}_{t_{1}}^{(\tau)}$ and $\bm{U}_{t_{2}}^{(\tau)}$ do not share components when $t_{1},t_{2}\in\mathcal{\tilde{N}}_{T}$. Thus, they are independent of each other due to Assumption 2.3, and $|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|$ for all $t\in\mathcal{\tilde{N}}_{T}$ are mutually independent. In turn, the events in the set
$\left\{\mathbb{I}\biggr{\{}|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|>\frac{1}{2}\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}\biggr{\}}:t\in\mathcal{\tilde{N}}_{T}\right\}$
are mutually independent. This fact allows us to use the Chernoff bound as follows:
$\mathbb{P}\Biggr{(}\sum_{t\in\mathcal{\tilde{N}}_{T}}\mathbb{I}\biggr{\{}|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|>\frac{1}{2}\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}\biggr{\}}\geq\frac{\alpha\tilde{N}_{T}}{2}\Biggr{)}\geq 1-\exp\left(-\frac{\alpha\tilde{N}_{T}}{8}\right),$
where $\alpha>0$ is defined in (47). Thus, we have
$\sum_{t\in\mathcal{\tilde{N}}_{T}}|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|$ $\geq\frac{\alpha\tilde{N}_{T}}{2}\cdot\frac{1}{2}\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}+\left(\tilde{N}_{T}-\frac{\alpha\tilde{N}_{T}}{2}\right)\cdot 0$
$=\frac{\alpha\tilde{N}_{T}}{4}\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}$ (48)
with probability at least $1-\exp\bigr{(}-\frac{\alpha\tilde{N}_{T}}{8}\bigr{)}$.
Finally, we have
$\sum_{t=\tau}^{T-1}\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau,i)}=0\}$ $\geq\sum_{t=\tau}^{T-1}\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau)}\equiv 0\}$
$=\sum_{t\in\mathcal{N}_{T}}\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}\geq\sum_{t\in\mathcal{\tilde{N}}_{T}}\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}.$
Combining this with (D) completes the proof. ∎
We have defined $\nu$ in the above lemma. We will show that the value of this constant is at least 1.
@@
Define $\nu:=\frac{\sqrt{M}\tau L_{\phi}\sigma_{u}}{\lambda}$. Then, $\nu\geq\Theta(1)$.
@@
For any $s\in\mathbb{S}^{M-1}$, we have
$|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|^{2}\leq\|s\|_{2}^{2}\cdot\|\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{2}=\|\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{2}\leq ML_{\phi}^{2}\|\bm{U}_{t}^{(\tau)}\|_{2}^{2},$ (49)
where the last inequality comes from (42). The expectation of the left-hand side of (49) is lower-bounded by $\lambda^{2}$ due to (45) and the right-hand side is upper-bounded by $\Theta(ML_{\phi}^{2}\tau^{2}\sigma_{u}^{2})$ due to (43). Thus, we have $\Theta(1)\leq\frac{ML_{\phi}^{2}\tau^{2}\sigma_{u}^{2}}{\lambda}$, which is equivalent to $\nu^{2}$. ∎
@@
Suppose that Assumptions 2.3 and 2.6 hold, and $\mathbb{P}(\bm{W}_{t}^{(\tau,i)}>0)=\mathbb{P}(\bm{W}_{t}^{(\tau,i)}<0)$ for all $t$ and $i$. Define $\nu:=\frac{\sqrt{M}\tau L_{\phi}\sigma_{u}}{\lambda}$. For a fixed $s\in\mathbb{S}^{M-1}$, it holds that
$\mathbb{P}\biggr{(}\sum_{t=\tau}^{T-1}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})>-\frac{\tilde{N}_{T}}{2}\cdot\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}\biggr{)}>1-\exp\biggr{(}-\Theta\biggr{(}\frac{\tilde{N}_{T}^{2}}{T\nu^{8}}\biggr{)}\biggr{)}.$ (50)
@@
Due to Assumption 2.6, we have
$|s_{i}\phi_{i}(\bm{U}_{t}^{(\tau)})|\leq|s_{i}|\cdot L_{\phi}(\|u_{t}\|_{2}+\cdots\|u_{t-\tau}\|_{2})$ (51)
for all $i=1,\dots,M$. Thus, we have
$|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|\leq L_{\phi}\sum_{i=1}^{M}|s_{i}|\cdot\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}\leq L_{\phi}\sqrt{M}\|s\|_{2}\cdot\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}=L_{\phi}\sqrt{M}\cdot\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}.$ (52)
Consider the constants $A_{1},\dots,A_{T}>0$. Then, we have
$\operatorname*{arg\,max}_{\begin{subarray}{c}|a_{t}|\leq A_{t},\\ t=\tau,\dots,T-1\end{subarray}}\mathbb{E}\Bigr{[}\exp\Bigr{(}\Bigr{(}\sum_{t=\tau}^{T-1}a_{t}\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})\Bigr{)}^{2}\Bigr{)}\Bigr{]}\subseteq\{\pm A_{1}\}\times\cdots\times\{\pm A_{T}\}$ (53)
since the function on the left-hand side is convex in $(a_{1},\dots,a_{T})$ and the maximum of the convex function is attained at extreme points. Due to (52), substituting $s^{T}\Phi(\bm{U}_{t}^{(\tau)})$ and $L_{\phi}\sqrt{M}\cdot\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}$ into $a_{t}$ and $A_{t}$ in (53) yields
$\left\|\sum_{t=\tau}^{T-1}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})\right\|_{\psi_{2}}\leq\left\|\sum_{t=\tau}^{T-1}L_{\phi}\sqrt{M}\cdot\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})\right\|_{\psi_{2}}$ (54)
since $\text{sgn}(\bm{W}_{t}^{(\tau,i)})$ has the same distribution as $-\text{sgn}(\bm{W}_{t}^{(\tau,i)})$ by the sign symmetry assumption.
Now, we analyze the right-hand side of (54). For simplicity, we let $\text{sgn}(\bm{W}_{t}^{(\tau,i)})=0$ for $t=0,\dots,\tau-1$ and also for $t=T,\dots,T+\tau-1$. Then, we have
$\sum_{t=\tau}^{T-1}\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})=\sum_{t=0}^{T-1}\left(\sum_{j=t}^{t+\tau}\text{sgn}(\bm{W}_{t}^{(\tau,i)})\right)\cdot\|u_{t}\|_{2}.$ (55)
@@
$\left\|\left(\sum_{j=t}^{t+\tau}\text{sgn}(\bm{W}_{t}^{(\tau,i)})\right)\cdot\|u_{t}\|_{2}\right\|_{\psi_{2}}\leq(\tau+1)\sigma_{u}$
due to Assumption 2.3. Given the filtration $\mathcal{F}^{i}=\bm{\sigma}\{\text{sgn}(\bm{W}_{t}^{(\tau,i)}):t=0,\dots,T-1\}$ and considering that $\mathbb{E}[\text{sgn}(\bm{W}_{t}^{(\tau,i)})]=0$, we can apply the property (28) to obtain
$\mathbb{E}\Biggr{[}\exp\Biggr{(}\theta\sum_{t=0}^{T-1}$ $\left(\sum_{j=t}^{t+\tau}\text{sgn}(\bm{W}_{t}^{(\tau,i)})\right)\cdot\|u_{t}\|_{2}\Biggr{)}\Biggr{]}$
$\leq\mathbb{E}\Biggr{[}\mathbb{E}\Biggr{[}\exp\Biggr{(}\theta\sum_{t=0}^{T-1}\left(\sum_{j=t}^{t+\tau}\text{sgn}(\bm{W}_{t}^{(\tau,i)})\right)\cdot\|u_{t}\|_{2}\Biggr{)}\Biggr{]}~{}\Biggr{|}~{}\mathcal{F}^{i}\Biggr{]}$
@@
for all $\theta\in\mathbb{R}$, which implies that the mean-zero variable (55) is sub-Gaussian and its norm is $\Theta(\sqrt{T}(\tau+1)\sigma_{u})$. In turn, due to (54), we arrive at
$\left\|\sum_{t=\tau}^{T-1}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})\right\|_{\psi_{2}}\leq\Theta(\sqrt{T}L_{\phi}\sqrt{M}(\tau+1)\sigma_{u}).$ (57)
@@
$\mathbb{P}\biggr{(}\sum_{t=\tau}^{T-1}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\cdot$ $\text{sgn}(\bm{W}_{t}^{(\tau,i)})>-\frac{\tilde{N}_{T}}{2}\cdot\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}\biggr{)}$
$\geq 1-\exp\biggr{(}-\Theta\biggr{(}\frac{\tilde{N}_{T}^{2}\lambda^{2}}{(\sqrt{T}L_{\phi}\sqrt{M}(\tau+1)\sigma_{u})^{2}\nu^{6}}\biggr{)}\biggr{)}$
$=1-\exp\biggr{(}-\Theta\biggr{(}\frac{\tilde{N}_{T}^{2}}{T\nu^{8}}\biggr{)}\biggr{)},$
which completes the proof. ∎
Corollary D.5.
Suppose that Assumptions 2.3 and 2.6 hold, and $\mathbb{P}(\bm{W}_{t}^{(\tau,i)}>0)=\mathbb{P}(\bm{W}_{t}^{(\tau,i)}<0)$ for all $t$ and $i$. Define $\nu:=\frac{\sqrt{M}\tau L_{\phi}\sigma_{u}}{\lambda}$ and
$z_{t}^{i}(s):=\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau,i)}=0\}$ $+s^{T}\Phi(\bm{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)}).$ (58)
Then, we have
$\mathbb{P}\biggr{(}\sum_{t=\tau}^{T-1}z_{t}^{i}(s)>\frac{\tilde{N}_{T}}{2}\cdot\Theta\Bigr{(}\frac{\lambda}{\nu^{3}}\Bigr{)}\biggr{)}\geq 1-\exp(-\Theta(\tilde{N}_{T}))-\exp\biggr{(}-\Theta\biggr{(}\frac{\tilde{N}_{T}^{2}}{T\nu^{8}}\biggr{)}\biggr{)}$ (59)
for any fixed $s\in\mathbb{S}^{M-1}$.
@@
One can apply the union bound on Lemma D.2 and Lemma D.4 to derive the conclusion. ∎
Lemma D.6.
Suppose that Assumptions 2.3 and 2.6 hold. Given $\delta\in(0,1]$, when $T\geq\Theta(\log(1/\delta))$, the inequality
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)-\sum_{t=\tau}^{T-1}z_{t}^{i}(\tilde{s})\geq-\Theta(T\|s-\tilde{s}\|_{2}L_{\phi}\sqrt{M}\tau\sigma_{u})$
holds for every $s,\tilde{s}\in\mathbb{S}^{M-1}$ with probability at least $1-\frac{\delta}{2}$.
Proof.
For $s,\tilde{s}\in\mathbb{S}^{M-1}$, we have
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)-\sum_{t=\tau}^{T-1}z_{t}^{i}(\tilde{s})$ $\geq-\sum_{t=\tau}^{T-1}|(s-\tilde{s})^{T}\Phi(\bm{U}_{t}^{(\tau)})|$
$\geq-\sum_{t=\tau}^{T-1}\|s-\tilde{s}\|_{2}\cdot L_{\phi}\sqrt{M}\cdot\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}$
$\geq-\sum_{t=0}^{T-1}\|s-\tilde{s}\|_{2}L_{\phi}\sqrt{M}(\tau+1)\|u_{t}\|_{2},$ (60)
where the first inequality is due to the triangle inequality and the second comes from (52). Assumption 2.3 tells that $\left\|\|u_{t}\|_{2}\right\|=\sigma_{u}$ and thus, we have
$\left\|\|s-\tilde{s}\|_{2}L_{\phi}\sqrt{M}(\tau+1)(\|u_{t}\|_{2}-\mathbb{E}[\|u_{t}\|_{2}])\right\|_{\psi_{2}}\leq\|s-\tilde{s}\|_{2}L_{\phi}\sqrt{M}(\tau+1)\cdot\Theta(\sigma_{u})$
@@
$\left\|\sum_{t=0}^{T-1}\|s-\tilde{s}\|_{2}L_{\phi}\sqrt{M}(\tau+1)(\|u_{t}\|_{2}-\mathbb{E}[\|u_{t}\|_{2}])\right\|_{\psi_{2}}\leq\|s-\tilde{s}\|_{2}L_{\phi}\sqrt{M}(\tau+1)\cdot\Theta(\sqrt{T}\sigma_{u}).$

### v2

@@ Appendix B Preliminaries on sub-Gaussian variables
$\mathbb{E}\left[\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}\right]=\Omega\Bigr(\frac{\lambda^{4}}{(\sqrt{M\tau}L_{\phi}\sigma_{u})^{3}}\Bigr).$ (40)
Now, recall the relationship
$\{\bm{W}_{t}^{(\tau)}=0\}\supseteq\{w_{t-1}=0,\dots,w_{t-\tau}=0\}\supseteq\{\xi_{t-1}=0,\dots,\xi_{t-\tau}=0\},$
which follows from Assumption 2.8. From the independence of $\xi_{i}$’s, we also have
$\mathbb{P}(\xi_{t-1}=0,\dots,\xi_{t-\tau}=0)=(1-p)^{\tau}>0.5,$
since $p<\frac{1}{2\tau}$. Then, one can write
$\mathbb{E}\left[\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}\right]$ $\geq\mathbb{E}\left[\mathbb{I}_{\pm}\{\xi_{t-1}=0,\dots,\xi_{t-\tau}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}\right]$ (41)
$=\mathbb{E}\left[\mathbb{I}_{\pm}\{\xi_{t-1}=0,\dots,\xi_{t-\tau}=0\}\right]\cdot\mathbb{E}\left[\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}\right]$
$\geq((1-p)^{\tau}-(1-(1-p)^{\tau})\cdot\mathbb{E}\left[\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}\right]$
$=(2(1-p)^{\tau}-1)\cdot\Omega\Bigr(\frac{\lambda^{4}}{(\sqrt{M\tau}L_{\phi}\sigma_{u})^{3}}\Bigr).$ (42)
We finally note that the term given in (42) is indeed positive since $(1-p)^{\tau}>0.5$. Using the definition of $\nu$ completes the proof. ∎
We have defined $\nu$ in the above lemma. We will show that the value of $\nu$ is bounded below by a positive constant.
@@
Define $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$. Then, $\nu=\Omega(1)$.
@@
For any $Z\in\mathbb{R}^{r\times M}$ such that $\|Z\|_{F}=1$, we have
$\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{2}\leq\|Z\|_{F}^{2}\cdot\|\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{2}=\|\Phi(\bm{U}_{t}^{(\tau)})\|_{2}^{2}\leq ML_{\phi}^{2}\|\bm{U}_{t}^{(\tau)}\|_{2}^{2},$ (43)
where the last inequality comes from (36). The expectation of the left-hand side of (43) is lower-bounded by $\lambda^{2}$ due to (39). Noting that the expectation and the $\psi_{2}$-norm of a nonnegative variable have the same order (see (24a)), the expectation of the right-hand side is upper-bounded by $O(ML_{\phi}^{2}\tau\sigma_{u}^{2})$ due to Lemma B.5. Thus, we have $\frac{ML_{\phi}^{2}\tau\sigma_{u}^{2}}{\lambda}=\Omega(1)$; in other words, $\nu^{2}=\Omega(1)$. This completes the proof. ∎
Now, we provide the crucial lemma to ensure that for a fixed $Z$, the term $\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}$ is positive with probability at least $1-\delta$.
@@
Suppose that Assumptions 2.4 and 2.5 hold. Define $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$. Given $\delta\in(0,1]$, when
$T=\Omega\left(\frac{\tau\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\log\Bigr(\frac{1}{\delta}\Bigr)\right),$ (44)
we have
$\sum_{t=\tau}^{T-1}\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}=\Omega\left(\frac{(2(1-p)^{\tau}-1)\cdot\lambda T}{2\nu^{3}}\right)$ (45)
for a fixed $Z\in\mathbb{R}^{r\times M}$ such that $\|Z\|_{F}=1$.
@@
Similar to (36) and (D), we have
$\sum_{t=\tau}^{T-1}\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}$ $\leq\sqrt{M}L_{\phi}\sum_{t=\tau}^{T-1}\sqrt{\sum_{i=0}^{\tau}\|u_{t-i}\|_{2}^{2}}\leq\sqrt{M}L_{\phi}\sum_{t=\tau}^{T-1}\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}.$ (46)
Now, we define a Bernoulli variable $\bm{\Xi}_{t}^{(\tau)}$ such that $\{\bm{\Xi}_{t}^{(\tau)}=0\}\Leftrightarrow\{\xi_{t-1}=0,\dots,\xi_{t-\tau}=0\}$. From (41), we know that $\{\bm{W}_{t}^{(\tau)}=0\}\supseteq\{\bm{\Xi}_{t}^{(\tau)}=0\}$. Thus, it suffices to prove the desired result with $\bm{\Xi}_{t}^{(\tau)}$ in place of $\bm{W}_{t}^{(\tau)}$.
Consider the constants $A_{1},\dots,A_{T}>0$. Then, for all $\theta\in\mathbb{R}$, we have
$\operatorname*{arg\,max}_{\begin{subarray}{c}|a_{t}|\leq A_{t},\\ t=\tau,\dots,T-1\end{subarray}}\mathbb{E}\Bigr[\exp\Bigr(\theta\Bigr(\sum_{t=\tau}^{T-1}a_{t}\cdot\bigr(\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}-\mathbb{E}[\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}]\bigr)\Bigr)^{2}\Bigr)\Bigr]$
$\hskip 227.62204pt\subseteq\{\pm A_{1}\}\times\cdots\times\{\pm A_{T}\},$ (47)
since the function on the left-hand side is convex in $(a_{1},\dots,a_{T})$ and the maximum of the convex function is attained at extreme points. Due to (46), substituting $\sum_{t=\tau}^{T-1}\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}$ into $a_{t}$ and $\sqrt{M}L_{\phi}(\tau+1)\sum_{t=\tau}^{T-1}\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}$ into $A_{t}$ in (D) yields
$\left\|\sum_{t=\tau}^{T-1}\|Z\Phi(\bm{U}_{t}^{(\tau)})\|_{2}\cdot\left(\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}-\mathbb{E}[\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}]\right)\right\|_{\psi_{2}}$
$\hskip 28.45274pt\leq\left\|\sqrt{M}L_{\phi}(\tau+1)\sum_{t=\tau}^{T-1}\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}\cdot\left(\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}-\mathbb{E}[\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}]\right)\right\|_{\psi_{2}}$ (48)
considering that $\bm{\Xi}_{t}^{(\tau)}$ is independent of any other variables and the expectation of $\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}-\mathbb{E}[\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}]$ is zero, in which case the sub-Gaussian norm can be determined by (24c).
Now, we analyze the right-hand side of (D). For simplicity, we define
$\Xi_{t}:=\begin{cases}0,&t=0,\dots,\tau-1,\\[6.0pt] \mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}-\mathbb{E}\bigr[\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}\bigr],&t=\tau,\dots,T-1,\\[6.0pt] 0,&t=T,\dots,T+\tau-1.\end{cases}$
Then, we have
$\sum_{t=\tau}^{T-1}\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}\cdot\Xi_{t}=\sum_{t=0}^{T-1}\Biggr(\sum_{j=t}^{t+\tau}\Xi_{j}\Biggr)\cdot\|u_{t}\|_{2}.$ (49)
@@
$\left\|\sum_{t=0}^{T-1}\Biggr(\sum_{j=t}^{t+\tau}\Xi_{j}\Biggr)\cdot\|u_{t}\|_{2}\right\|_{\psi_{2}}\leq(\tau+1)\sigma_{u}$
due to Assumption 2.5. Given the filtration $\mathcal{F}^{i}=\bm{\sigma}\{\Xi_{t}:t=0,\dots,T-1\}$ and considering that $\mathbb{E}[\Xi_{t}]=0$, we can apply the property (24c) to obtain
$\mathbb{E}\Biggr[\exp\Biggr(\theta\sum_{t=0}^{T-1}\Biggr(\sum_{j=t}^{t+\tau}\Xi_{j}\Biggr)\cdot\|u_{t}\|_{2}\Biggr)\Biggr]$ $\leq\mathbb{E}\Biggr[\mathbb{E}\Biggr[\exp\Biggr(\theta\sum_{t=0}^{T-1}\Biggr(\sum_{j=t}^{t+\tau}\Xi_{j}\Biggr)\cdot\|u_{t}\|_{2}\Biggr)\Biggr]~\Biggr|~\mathcal{F}^{i}\Biggr]$
@@
for all $\theta\in\mathbb{R}$, which implies that the mean-zero variable (49) is sub-Gaussian and its norm is $O(\sqrt{T}(\tau+1)\sigma_{u})$. In turn, due to (D), we arrive at
$\left\|\sum_{t=\tau}^{T-1}Z\Phi(\bm{U}_{t}^{(\tau)})\cdot\Xi_{t}\right\|_{\psi_{2}}=O(\sqrt{TM}L_{\phi}(\tau+1)\sigma_{u}).$ (51)
@@
$\mathbb{P}\biggr(\sum_{t=\tau}^{T-1}Z\Phi(\bm{U}_{t}^{(\tau)})\cdot\Xi_{t}$ $>-\Omega\left(\frac{(2(1-p)^{\tau}-1)\cdot\lambda T}{2\nu^{3}}\right)\biggr)$
$\geq 1-\exp\biggr(-\Omega\biggr(\frac{(2(1-p)^{\tau}-1)^{2}\lambda^{2}T^{2}}{(\sqrt{TM}L_{\phi}(\tau+1)\sigma_{u})^{2}\nu^{6}}\biggr)\biggr)$
$=1-\exp\biggr(-\Omega\biggr(\frac{(2(1-p)^{\tau}-1)^{2}\cdot T}{\tau\nu^{8}}\biggr)\biggr).$
We derive from (42) that
$\mathbb{E}\left[\sum_{t=\tau}^{T-1}Z\Phi(\bm{U}_{t}^{(\tau)})\cdot\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}\right]=\Omega\Bigr(\frac{(2(1-p)^{\tau}-1)\cdot\lambda T}{\nu^{3}}\Bigr).$
Since $\Xi_{t}=\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}-\mathbb{E}\bigr[\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}\bigr]$, we arrive at
$\mathbb{P}\biggr(\sum_{t=\tau}^{T-1}Z\Phi(\bm{U}_{t}^{(\tau)})\cdot\mathbb{I}_{\pm}\{\bm{\Xi}_{t}^{(\tau)}=0\}$ $>\Omega\left(\frac{(2(1-p)^{\tau}-1)\cdot\lambda T}{2\nu^{3}}\right)\biggr)$
$\hskip 14.22636pt\geq 1-\exp\biggr(-\Omega\biggr(\frac{(2(1-p)^{\tau}-1)^{2}\cdot T}{\tau\nu^{8}}\biggr)\biggr).$ (52)
Since we have $\{\bm{W}_{t}^{(\tau)}=0\}\supseteq\{\bm{\Xi}_{t}^{(\tau)}=0\}$, establishing a lower bound of $1-\delta$ on the right-hand side of (52) suffices to conclude the proof. ∎
We now study the effect of perturbing $Z\in\mathbb{R}^{r\times M}$.
Lemma D.6.
Suppose that Assumptions 2.4 and 2.5 hold. Given $\delta\in(0,1]$, when $T=\Omega(\log(2/\delta))$, the inequality
$\sum_{t=\tau}^{T-1}\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}-\sum_{t=\tau}^{T-1}\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|\tilde{Z}\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}\geq-O(T\|Z-\tilde{Z}\|_{F}L_{\phi}\sqrt{M}\tau\sigma_{u})$
holds for every $Z,\tilde{Z}\in\mathbb{R}^{r\times M}$ with probability at least $1-\frac{\delta}{2}$.
@@
For simplicity, we define $\bar{f}_{t}(Z):=\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}$. For $Z,\tilde{Z}\in\mathbb{R}^{r\times M}$, we have
[…]
$\geq-\sum_{t=\tau}^{T-1}\|Z-\tilde{Z}\|_{F}\cdot L_{\phi}\sqrt{M}\cdot\sum_{j=0}^{\tau}\|u_{t-j}\|_{2}$
$\geq-\sum_{t=0}^{T-1}\|Z-\tilde{Z}\|_{F}L_{\phi}\sqrt{M}(\tau+1)\|u_{t}\|_{2},$ (53)
where the first inequality is due to the triangle inequality and the second comes from (46). Assumption 2.5 tells that $\left\|\|u_{t}\|_{2}\right\|\leq\sigma_{u}$ and thus, we have
$\left\|\|Z-\tilde{Z}\|_{F}L_{\phi}\sqrt{M}(\tau+1)(\|u_{t}\|_{2}-\mathbb{E}[\|u_{t}\|_{2}])\right\|_{\psi_{2}}=\|Z-\tilde{Z}\|_{F}L_{\phi}\sqrt{M}(\tau+1)\cdot O(\sigma_{u})$
@@
$\left\|\sum_{t=0}^{T-1}\|Z-\tilde{Z}\|_{F}L_{\phi}\sqrt{M}(\tau+1)(\|u_{t}\|_{2}-\mathbb{E}[\|u_{t}\|_{2}])\right\|_{\psi_{2}}=\|Z-\tilde{Z}\|_{F}L_{\phi}\sqrt{M}(\tau+1)\cdot O(\sqrt{T}\sigma_{u}).$